# Goal-Oriented Status Updating for Real-time Remote Inference over Networks with Two-Way Delay

### v3

@@ II-A Communication Network Model
Consider that the communication network may have multiple routes $\{1,2,\ldots,N\}$ connecting the transmitter and the receiver. The network directs each submitted packet to one of these available routes. If a packet is transmitted through a specific route, its corresponding feedback is also transmitted through that same route.
Let the $i$-th packet transmission epoch consist of the time slots $A_{i-1},A_{i-1}+1,\ldots,A_{i}-1$. Define $T_{i}(l_{i})=D_{i}-S_{i}$ as the transmission delay incurred in the $i$-th epoch and $F_{i}=A_{i}-D_{i}$ as the feedback delay. Additionally, let $C_{i}$ represent the delay state in the $i$-th epoch, with $c_{i}\in\{1,2,\ldots,N\}$ denoting its realization. The delay state $C_{i}$ is determined by the route used in the $i$-th packet transmission epoch. Each route has distinct transmission and feedback delay distributions, and the transmission delay also varies with the packet length. The delay is at least one time slot for each packet or ACK transmission. We assume that $C_{i}$ evolves according to a finite-state ergodic Markov chain with transition probabilities $p_{ij}$, where $i,j\in\{1,2,\ldots,N\}$, and $p_{ij}$ represents the probability of transitioning from state $i$ to state $j$. The Markov chain makes a single transition at time slot $A_{i}$ and none otherwise.
The receiver detects the delay state $C_{i}$ at the packet delivery time $D_{i}$ and embeds this information into the ACK message. The ACK notifies the transmitter of (i) the delay state experienced during the previous packet transmission, (ii) the completion of the previous packet transmission, and (iii) that the next packet can be submitted to the network. The network is reliable, meaning no packet is lost during transmission.
@@ II-B Inference Error as a Function of AoI and Packet Length
These trained neural networks are used to predict the real-time value of the target signal $Y_{t}$ on the receiver side. We assume that the process ${\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}}$ is stationary, and that the processes ${\{\big{(}Y_{t},V_{t}\big{)},t=0,1,\ldots\}}$ and ${\{\big{(}\Delta(t),d(t)\big{)},t=0,1,\ldots\}}$ are independent. The average inference error at time slot $t$, given $\Delta(t)=\delta$ and $d(t)=l$, is expressed as:
@@ III-A Co-design Problem Formulation
Let $\Pi_{l}$ represent the set of all causal scheduling policies $\pi=(f,\ell,g)$ with a time-invariant packet length sequence =${\ell=\{l,l,\ldots\}}$ for $l\in\{1,2,\ldots,B\}$:
@@
Problem (6)-(7) is a more general formulation than that considered in [17, Section IV-A], as it accounts for both (i) the two-way delay between the transmitter and receiver, and (ii) the Markovian, time-varying nature of the two-way delay distribution in multi-route communication networks.
@@ III-B Optimal Solution to (6)-(7)
where $\delta\in\mathbb{Z}^{+}$, $c_{i}\in\{1,2,\ldots,N\}$, and $h_{1}(\delta,c_{i})$ is the relative value function. We need to solve the Bellman optimality equation (III-B) to find an optimal solution to (6). This involves determining the optimal buffer position $b_{i+1}^{*}$ and waiting time $\tau_{i+1}^{*}$ for each state $\big{(}\delta,c_{i}\big{)}$. By exploiting the structural properties of the Bellman optimality equation (III-B), we are able to obtain a closed-form solution to (6), as asserted in the following theorem.
@@
Step 2: Since the delay state $c_{i}$ in the $i$-th epoch is a sufficient statistic for determining $b_{i+1}^{*}$ for all $i$, we can construct a new SMDP for determining $b_{i+1}^{*}$ using $c_{i}$ as the state. In the new SMDP, we fix the waiting time decisions by using optimal threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$. Moreover, in the new SMDP, the decision time is $D_{i+1}$ instead of $A_{i}$. The Bellman optimality equation of this new SMDP is expressed as:
@@
The optimal scheduling policy for the inner optimization problem (6), as outlined in Theorem 1, is well-structured. Each optimal waiting time $\tau_{i+1}^{*}$ is determined by an index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$. The index function $\gamma\big{(}\cdot\big{)}$, defined in equation (9), can be readily computed for any state $\big{(}\delta,c_{i}\big{)}$ using the conditional distribution of $T_{i+1}(l)$ given $C_{i}=c_{i}$. Although the index-based threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$ closely resembles the waiting time rule provided in [17, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state $C_{i}=c_{i}$. Because problem (6)-(7) involves non-zero feedback delay and time-varying delay distribution, the existence of such a simple waiting time rule was not evident and required additional technical efforts for its derivation.
Furthermore, according to equation (1), the optimal buffer position $b_{i+1}^{*}$ is designed by minimizing the relative inference error in the interval between $D_{i+1}$ and $D_{i+2}$, given that the waiting time is determined by the optimal rule in (10). Each $b_{i+1}^{*}$ depends solely on the delay state $C_{i}=c_{i}$ in the $i$-th epoch and is independent of the AoI state $\delta=\Delta(A_{i})$. For the case of i.i.d. transmission delay and immediate feedback, the optimal buffer position sequence is time-invariant, as shown in [17, Theorem 1]. However, when the feedback delay is non-zero and the delay state $C_{i}$ changes according to a Markov chain, achieving optimal performance requires that the optimal buffer position $b_{i+1}^{*}$ is a function of the delay state $C_{i}=c_{i}$, and this function is exactly given by (1).
@@ IV-B Optimal Solution to (14)
[…]
@@
The Bellman optimality equations (IV-B) and (2) for the SMDP (14) can be solved using policy iteration, value iteration, or linear programming [23, Chapter 11.4.4]. In particular, policy iteration is efficient for solving SMDPs [17, 23]. The time complexity for solving the simplified Bellman equation (2) is significantly lower than that for solving (IV-B): The policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2). The policy evaluation step is identical for both equations.
@@ V-A Model-Based Evaluation
Higher-order auto-regressive (AR) linear time-invariant systems are widely used models for various practical scenarios, such as control systems incorporating memory and delay [43, 44, 45], and wireless communication channels [46]. Motivated by this, we consider a remote inference problem for an AR process, where the target signal evolves according to:
@@
where the noise $W_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{W}^{2}$ and $Y_{t}\in\mathbb{R}$. Let $V_{t}=Y_{t}+N_{t}$ be the noisy observation of the target $Y_{t}$, where $N_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{N}^{2}$. We set $\sigma_{W}^{2}=0.01$, $\sigma_{N}^{2}=0.001$ and construct an AR($10$) process with coefficients $a_{2}=0.05$ and $a_{10}=0.9$. The remaining coefficients of the AR($10$) process are zero. We consider a quadratic loss function $L_{2}(y,\hat{y})=\lVert y-\hat{y}\rVert_{2}^{2}$. In this experiment, the goal is to infer the target signal $Y_{t}$ by using the data packet $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$. A linear MMSE estimator achieves the optimal inference error since $X_{t-\delta}^{l}$ and $Y_{t}$ are jointly Gaussian, and the loss function is quadratic. Fig. 3 shows the inference error for AoI values ${\delta=1,2,\ldots,50}$ and packet lengths $l=1,2,4,6,8,10$.
Using this AR process, we illustrate the performance of the optimal policy derived for the scheduling problem specified in (6)-(7) under time-invariant packet length selection.
Let the number of delay states be $N=2$, implying that the delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $p_{11}$, $p_{12}$, $p_{21}$, and $p_{22}$. During the simulations, we assume $p_{12}=p_{21}$, ensuring that each delay state has a steady-state probability of $\frac{1}{2}$. The transmission delay $T_{i}(l)$ for each delay state is given by
@@
where $\sigma$ is a scaling parameter. Additionally, the feedback delay $F_{i}$ is $1$ time slot when $c_{i}=1$ and $3$ time slots when $c_{i}=2$. The maximum allowed packet length is $10$, and the buffer size is $B=75$.
We control the delay memory by varying the sum of transition probabilities $\alpha=p_{12}+p_{21}$. When $\alpha=1$, i.e., ${p_{12}=p_{21}=\frac{1}{2}}$, both delay states are equally likely in the $i$-th epoch, independent of the delay state in the $(i-1)$-th epoch. Consequently, the delay distribution is i.i.d.. However, as $\alpha$ deviates from $1$, the delay memory increases, and the delay state in the $(i-1)$-th epoch becomes more informative about the delay state in the $i$-th epoch.
Fig. 4 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 00 to $2$ with $\alpha=\frac{1}{20}$:
@@
Fig. (5) presents the time-average inference error achieved by the following two scheduling policies as the parameter $\alpha$ varies from 00 to $2$ with $\sigma=\frac{5}{2}$:
@@
    Policy outlined in [17, Theorem 1] with $l=5$. This policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
@@ V-B Trace-Driven Evaluation
Ensuring the safe operation of robotic systems often requires evaluating their state based on observations from remote sources. This makes the prediction problem critical for monitoring the safety of such systems. We consider a specific scenario, illustrated in Fig. 6, where a camera remotely monitors a control system, e.g., cart-pole. The camera captures video frames and transmits them to a receiver over a network with distinct delay states. Due to the high dimensionality of the video data, transmitting the sequence of frames takes multiple time slots. As a result, the most recently delivered sequence of video frames is generated $\delta$ time slots ago. Upon reception, the frames undergo data pre-processing to extract the cart position and the pole angle information, ${X_{t-\delta}=\big{(}(\psi_{t-\delta},Z_{t-\delta}),\ldots,(\psi_{t-\delta-l+1},Z_{t-\delta-l+1})\big{)}}$. Subsequently, a predictor uses the most recent available information to estimate the current pole angle and cart position ${Y_{t}=(\psi_{t},Z_{t})}$.
Figure 7: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and packet lengths ${l=1,2,4,6,8,10}$.
@@
We consider two delay states $N=2$. The delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $p_{11}$, $p_{12}$, $p_{21}$, and $p_{22}$. We set $p_{12}=p_{21}$ to ensure equal steady-state probabilities for both delay states. The transmission delay $T_{i}(l)$ for each delay state is given by
$T_{i}(l)=\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ 5+\lceil 5\sigma l\rceil,&\text{if }c_{i}=2,\end{cases}$ (20)
where $\sigma$ is a scaling parameter. The feedback delay $F_{i}$ is identical to that in the AR process experiment. We set the sum of transition probabilities as $\alpha=p_{12}+p_{21}=\frac{1}{20}$.
(a) $\sigma=\frac{1}{2}$
(b) $\sigma=\frac{5}{2}$
Figure 8: Time-average inference error vs. the buffer size B for ${\sigma=\frac{1}{2}}$ and ${\sigma=\frac{5}{2}}$.
@@
Fig. 8(a) and Fig. 8(b) show the time-average inference error achieved by the optimal policy in Theorem 2 for buffer size values $B$, ranging from $1$ to $10$, with ${\sigma=\frac{1}{2}}$ and ${\sigma=\frac{5}{2}}$, respectively. As the buffer size increases, the performance improvement diminishes and becomes less significant. Therefore, we fix the buffer size $B$ at $10$ for this experiment.
Fig. 9 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 00 to $3$:
@@
    Generate-at-will, zero-wait, $l=5$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(5,5,\ldots)$.
The result in Fig. 9 highlights the importance of adjusting the packet length over time in response to varying delay conditions. The optimal policy in Theorem 2 outperforms the other two policies, which use constant packet length, achieving up to $25\%$ reduction in average inference error.

### v4

@@ II-A Communication Network Model
Next-generation communication networks are expected to accommodate multiple routes between a transmitter–receiver pair through a variety of intermediate network components, such as LEO satellites, inter-satellite links, and ground gateways. The experienced delay may therefore vary significantly depending on the specific components traversed by a packet and the manner in which it is handled during transmission. For example, when a packet is delivered entirely via inter-satellite links under moderate network load, the dominant delay component is propagation delay. In contrast, when inter-satellite link unavailability requires routing through intermediate ground stations, or when packets undergo store-and-forward transmission over long distances in LEO satellites under heavy network load, the resulting delay may involve either substantially larger propagation components or considerable store-and-forward waiting times. In such networks, the receiver can characterize the realized end-to-end delay behavior through a detection rule that is consistent with its observation capabilities (e.g., based on measured packet arrival times). The precision of this characterization depends on the receiver’s observation capability.
Furthermore, as in any practical communication system, physical-layer transmission errors and link-layer retransmissions may occur along these routes. In this work, we adopt a high-level network-layer perspective, where such effects are incorporated into the end-to-end transmission delay. Specifically, the delay experienced by a packet accounts for processing and propagation delays, as well as potential local retransmissions and error recovery mechanisms within the network. This modeling choice allows us to focus on the scheduler’s decision-making at the source, rather than on lower-layer protocol operations.
@@
Let the $i$-th packet transmission epoch consist of the time slots $A_{i-1},A_{i-1}+1,\ldots,A_{i}-1$. Define $T_{i}(l_{i})=D_{i}-S_{i}$ as the transmission delay incurred in the $i$-th epoch and $F_{i}=A_{i}-D_{i}$ as the feedback delay. In addition, let $C_{i}$ denote the network delay state in the $i$-th epoch, with its realization $c_{i}$. Fig. 2 provides an illustration of the $i$-th and the $(i+1)$-th packet transmission epochs.
The network delay state $C_{i}$ in the $i$-th epoch can take values in $\{1,2,\ldots,N\}$, each representing a distinct delay regime that the network may experience during the transmission of the $i$-th data packet and that can be distinguished by the receiver. Each delay regime is associated with different transmission and feedback delay distributions, and the transmission delay also varies with the packet length. The delay is at least one time slot for each packet or ACK transmission.
We assume that $C_{i}$ evolves according to a finite-state ergodic Markov chain with transition probabilities $p_{ij}$, where $i,j\in\{1,2,\ldots,N\}$ and $p_{ij}$ represents the probability of transitioning from state $i$ to state $j$. The Markov chain makes a single transition at time slot $A_{i}$ and none otherwise.
The receiver detects the network delay state $C_{i}$ at the packet delivery time $D_{i}$ and embeds this information into the ACK message. The ACK notifies the transmitter of (i) the delay state experienced during the previous packet transmission, (ii) the completion of the previous packet transmission, and (iii) that the next packet can be submitted to the network. The network is reliable, meaning no packet is lost during transmission.
@@ II-B Inference Error as a Function of AoI and Packet Length
These pre-trained neural networks are used at the destination to predict the real-time value of the target signal $Y_{t}$ based on the most recently received data packet. We assume that the stochastic process $\{\big(Y_{t},V_{t}\big),t=0,1,\ldots\}$} is stationary, implying that the statistical relationship between the source signal and the target signal is time-invariant. In addition, we further assume that the process $\{\big(Y_{t},V_{t}\big),t=0,1,\ldots\}$} is independent of the process $\{\big(\Delta(t),d(t)\big),t=0,1,\ldots\}$}. Accordingly, the average inference error at time slot $t$, conditioned on $\Delta(t)=\delta$ and $d(t)=l$, is expressed as:
@@ III-A Co-design Problem Formulation
Let $\Pi_{l}$ represent the set of all causal scheduling policies $\pi=(f,\ell,g)$ with a time-invariant packet length sequence $\ell=\{l,l,\ldots\}$} for $l\in\{1,2,\ldots,B\}$:
@@
Problem (6)-(7) is a more general formulation than that considered in [8, Section IV-A], as it accounts for both (i) the two-way delay between the transmitter and receiver, and (ii) the Markovian, time-varying nature of the two-way delay distribution in next-generation communication networks.
@@ III-B Optimal Solution to (6)-(7)
where $\delta\in\mathbb{Z}^{+}$, $c_{i}\in\{1,2,\ldots,N\}$, and $h_{1}(\delta,c_{i})$ is the relative value function.
We need to solve the Bellman optimality equation (III-B) to find an optimal solution to (6). This involves determining the optimal buffer position $b_{i+1}^{*}$ and waiting time $\tau_{i+1}^{*}$ for each state $\big(\delta,c_{i}\big)$. By exploiting the structural properties of the Bellman optimality equation (III-B), we are able to obtain a closed-form solution to (6), as asserted in the following theorem.
@@
Step 2: Since the delay state $c_{i}$ in the $i$-th epoch is a sufficient statistic for determining $b_{i+1}^{*}$ for all $i$, we can construct a new SMDP for determining $b_{i+1}^{*}$ using $c_{i}$ as the state. In the new SMDP, we fix the waiting time decisions by using the optimal threshold rule $\tau_{l}\big(\delta,c_{i}\big)$. Moreover, in the new SMDP, the decision time is $D_{i+1}$ instead of $A_{i}$. The Bellman optimality equation of this new SMDP is expressed as:
@@
The optimal scheduling policy for the inner optimization problem (6), as outlined in Theorem 1, is well-structured. Each optimal waiting time $\tau_{i+1}^{*}$ is determined by an index-based threshold rule $\tau_{l}\big(\delta,c_{i}\big)$. The index function $\gamma\big(\cdot\big)$, defined in equation (9), can be readily computed for any state $\big(\delta,c_{i}\big)$ using the conditional distribution of $T_{i+1}(l)$ given $C_{i}=c_{i}$. Although the index-based threshold rule $\tau_{l}\big(\delta,c_{i}\big)$ closely resembles the waiting time rule provided in [8, Theorem 1], it is not solely dependent on the AoI but also incorporates the delay state $C_{i}=c_{i}$. Because problem (6)-(7) involves non-zero feedback delay and time-varying delay distribution, the existence of such a simple waiting time rule was not evident and required additional technical efforts for its derivation.
Furthermore, according to equation (1), the optimal buffer position $b_{i+1}^{*}$ is designed by minimizing the relative inference error in the interval between $D_{i+1}$ and $D_{i+2}$, given that the waiting time is determined by the optimal rule in (10). Each $b_{i+1}^{*}$ depends solely on the delay state $C_{i}=c_{i}$ in the $i$-th epoch and is independent of the AoI state $\delta=\Delta(A_{i})$.
For the case of i.i.d. transmission delay and immediate feedback, the optimal buffer position sequence is time-invariant, as shown in [8, Theorem 1]. However, when the feedback delay is non-zero and the delay state $C_{i}$ changes according to a Markov chain, achieving optimal performance requires that the optimal buffer position $b_{i+1}^{*}$ is a function of the delay state $C_{i}=c_{i}$, and this function is exactly given by (1).
@@ IV-B Optimal Solution to (14)
$h_{2}\big(\delta,d,c_{i}\big)=\min_{\begin{subarray}{c}l_{i+1}\in\{1,2,\ldots,B\}\end{subarray}}\Bigg\{$
@@
The Bellman optimality equations (IV-B) and (2) for the SMDP (14) can be solved using policy iteration, value iteration, or linear programming [23, Chapter 11.4.4]. In particular, policy iteration is efficient for solving SMDPs [23, 8]. The time complexity for solving the simplified Bellman equation (2) via policy iteration is significantly lower than that for solving (IV-B). Specifically, the policy improvement step for (IV-B) has a time complexity of $O(\tau_{\text{bound}}^{2}B^{3})$, while it is $O(\tau_{\text{bound}}^{2}B+\tau_{\text{bound}}B^{3}+\tau_{\text{bound}}B^{2})=O(\max\{\tau_{\text{bound}}^{2}B,\tau_{\text{bound}}B^{3}\})$ for the simplified equation (2). The policy evaluation step is identical for both equations.
@@ V-A Model-Based Evaluation
Higher-order auto-regressive (AR) linear time-invariant systems are widely used models for various practical scenarios, such as control systems incorporating memory and delay [43, 44, 45], and wireless communication channels [46]. Motivated by this, we consider a remote inference problem for an AR process, where the target signal $Y_{t}\in\mathbb{R}$ evolves according to:
@@
where the noise $W_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{W}^{2}$. Let $V_{t}=Y_{t}+N_{t}$ be the noisy observation of $Y_{t}$, where $N_{t}\in\mathbb{R}$ is zero-mean Gaussian with variance $\sigma_{N}^{2}$.
In this experiment, we set $\sigma_{W}^{2}=0.01$, $\sigma_{N}^{2}=0.001$ and construct an AR($10$) process with coefficients $a_{2}=0.05$ and $a_{10}=0.9$. The remaining coefficients of the AR($10$) process are zero. We consider a quadratic loss function $L_{2}(y,\hat{y})=\lVert y-\hat{y}\rVert_{2}^{2}$}. The goal is to infer the target signal $Y_{t}$ by using the data packet $X_{t-\delta}^{l}=(V_{t-\delta},V_{t-\delta-1},\ldots,V_{t-\delta-l+1})$. A linear MMSE estimator achieves the optimal inference error since $X_{t-\delta}^{l}$ and $Y_{t}$ are jointly Gaussian, and the loss function is quadratic. Fig. 3 shows the inference error for AoI values $\delta=1,2,\ldots,50$} and packet lengths $l=1,2,4,6,8,10$.
Using this AR process, we illustrate the performance of the optimal policy derived for the scheduling problem specified in (6)-(7) under time-invariant packet length selection. For this experiment, we set the maximum allowable packet length to $l=10$ and the buffer size to $B=75$.
Let the number of network delay states be $N=2$, implying that the delay state $C_{i}$ in the $i$-th epoch follows a two-state Markov chain with transition probabilities $\{p_{jk}\}_{j,k\in\{1,2\}}$. During the simulations, we assume $p_{12}=p_{21}$, ensuring that each delay state has a steady-state probability of $\frac{1}{2}$.
We assume a simple transmission delay model that scales with the packet length and the current network conditions. The transmission delay $T_{i}(l)$ for each delay state is given by
@@
where $\sigma$ is a scaling parameter. On the other hand, the corresponding feedback delay $F_{i}$ is given by
$F_{i}=\begin{cases}1,&\text{if }c_{i}=1,\\ 3,&\text{if }c_{i}=2.\end{cases}$ (20)
We control the delay memory by varying the sum of transition probabilities $\alpha=p_{12}+p_{21}$. When $\alpha=1$, i.e., $p_{12}=p_{21}=\frac{1}{2}$}, both delay states are equally likely in the $i$-th epoch, independent of the delay state in the (i1)(i-1)-th epoch. Consequently, the delay distribution is i.i.d.. However, as $\alpha$ deviates from $1$, the delay memory increases, and the delay state in the (i1)(i-1)-th epoch becomes more informative about the delay state in the $i$-th epoch.
Fig. 4 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 0 to $2$ with $\alpha=\frac{1}{20}$:
@@
Fig. (5) presents the time-average inference error achieved by the following two scheduling policies as the parameter $\alpha$ varies from 0 to $2$ with $\sigma=\frac{5}{2}$:
@@
    Policy outlined in [8, Theorem 1] with $l=5$. This policy assumes that both delay states are equally likely at each epoch, independent of the history, for any $\alpha$.
@@ V-B Trace-Driven Evaluation
Ensuring the safe operation of robotic systems often requires evaluating their state based on observations from remote sources. This makes the prediction problem critical for monitoring the safety of such systems. We consider a specific scenario, illustrated in Fig. 6, where a camera remotely monitors a control system, e.g., a cart-pole. The camera captures video frames and transmits them to a receiver over a network with distinct delay states. Due to the high dimensionality of the video data, transmitting the sequence of frames takes multiple time slots. As a result, the most recently delivered sequence of video frames is generated $\delta$ time slots ago. Upon reception, the frames undergo data pre-processing to extract the cart position and the pole angle information, $X_{t-\delta}=\big((\psi_{t-\delta},Z_{t-\delta}),\ldots,(\psi_{t-\delta-l+1},Z_{t-\delta-l+1})\big)$}. Subsequently, a predictor uses the most recent available information to estimate the current pole angle and cart position $Y_{t}=(\psi_{t},Z_{t})$}.
Figure 7: Inference error of the cart-pole experiment for AoI values $\delta=1,2,\ldots,80$ and packet lengths l=1,2,4,6,8,10{l=1,2,4,6,8,10}.
@@
We consider a slightly more diverse network environment in this experiment by assuming three network delay states, i.e., $N=3$. Accordingly, the delay state $C_{i}$ in the $i$-th epoch follows a three-state Markov chain with transition probabilities $\{p_{jk}\}_{j,k\in\{1,2,3\}}$. We set the self-transition probabilities to $p_{11}=p_{22}=p_{33}=1-\alpha$, and all remaining transition probabilities to $\frac{\alpha}{2}$. Due to the symmetry of the Markov chain, the steady-state probability of each delay state is $\frac{1}{3}$. Throughout the simulation, we set $\alpha=\frac{1}{50}$.
The transmission delay corresponding to each state has two components: (i) a constant $10$ time slot component, and (ii) a component that scales depending on the packet length and the network delay condition. For example, the first component may represent the unavoidable, relatively large propagation delay in non-terrestrial networks. The other part may represent the varying levels of store-and-forward operation of the satellites, as well as the increase in the transmission delay due to packet lengths. As a result, the transmission delay $T_{i}(l)$ for each delay state is given by
$T_{i}(l)=10+\begin{cases}\lceil\sigma l\rceil,&\text{if }c_{i}=1,\\ \lceil 2\sigma l\rceil,&\text{if }c_{i}=2,\\ \lceil 3\sigma l\rceil,&\text{if }c_{i}=3,\end{cases}$ (21)
where $\sigma$ is a scaling parameter. On the other hand, the feedback delay $F_{i}$ for each delay state is given by
$F_{i}=\begin{cases}1,&\text{if }c_{i}=1,\\ 2,&\text{if }c_{i}=2,\\ 3,&\text{if }c_{i}=3.\end{cases}$ (22)
(a) $\sigma=\frac{1}{5}$
(b) $\sigma=\frac{9}{5}$
Figure 8: Time-average inference error vs. the buffer size B for $\sigma=\frac{1}{5}$} and $\sigma=\frac{9}{5}$}.
@@
Fig. 8(a) and Fig. 8(b) show the time-average inference error achieved by the optimal policy in Theorem 2 for buffer size values $B$, ranging from $1$ to $10$, with $\sigma=\frac{1}{5}$} and $\sigma=\frac{9}{5}$}, respectively. As the buffer size increases, the performance improvement diminishes and becomes less significant. Therefore, we fix the buffer size $B$ at $10$ for this experiment.
Fig. 9 presents the time-average inference error achieved by the following three scheduling policies as the parameter $\sigma$ changes from 0 to $2$:
@@
    Generate-at-will, zero-wait, $l=3$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(3,3,\ldots)$.
@@
    Generate-at-will, zero-wait, $l=6$ policy: $(f,\ell,g)$ such that $f=g=(0,0,\ldots)$ and $\ell=(6,6,\ldots)$.
The result in Fig. 9 highlights the importance of adjusting the packet length over time in response to varying delay conditions. The optimal policy in Theorem 2 outperforms the other two policies, which use constant packet length, achieving up to $15.7\%$ reduction in average inference error.
TABLE I: Normalized time complexity comparison for solving the original Bellman optimality equation (Orig. BE) (IV-B) and the simplified Bellman optimality equation (Simp. BE) (2), both with and without the pre-computed $\gamma(\cdot)$ in (9), using the policy iteration algorithm, for buffer sizes $B\in\{1,5,10\}$.
Buffer size $B$
Bellman equation $B=1$ $B=5$ $B=10$
Orig. BE $\mathcal{T}_{1}$ $\mathcal{T}_{5}$ $\mathcal{T}_{10}$
Simp. BE $\approx\mathcal{T}_{1}$ $0.8556\mathcal{T}_{5}$ $0.4527\mathcal{T}_{10}$
Simp. BE + pre-computed $\gamma(\cdot)$ $\approx\mathcal{T}_{1}$ $0.8531\mathcal{T}_{5}$ $0.4464\mathcal{T}_{10}$
We finally present an empirical comparison of the normalized time complexity for solving the original Bellman optimality equation (IV-B) and the simplified Bellman optimality equation (2), provided in Section IV, using the policy iteration algorithm. For the simplified Bellman optimality equation (2), the time complexity is reported for two cases: when the index function $\gamma(\cdot)$ in (9) is pre-computed and when it is not initially available. The results are summarized in Table I for buffer sizes $B\in\{1,5,10\}$. Let $\mathcal{T}_{k}$ denote the time required to solve the original Bellman equation for buffer size $B=k$. The corresponding solution times for the simplified Bellman equation, both with and without pre-computed $\gamma(\cdot)$, are expressed relative to $\mathcal{T}_{k}$.
For a small buffer size of $B=1$, the time complexities of solving the two equations are nearly identical. This is because, when $B=1$, the action space is extremely limited: the only possible choice for the buffer position action $b_{i+1}$ at decision time $A_{i}$ is zero, and the only possible packet length action $l_{i+1}$ is one. Consequently, the reduction in complexity of the policy improvement step of the policy iteration algorithm is not observable. In contrast, as the buffer size increases, solving the simplified Bellman equation has notably lower time complexity than solving the original Bellman optimality equation, achieving more than a two-fold reduction at buffer size $B=10$. Moreover, the computation of the index function $\gamma(\cdot)$, which is required for solving the simplified Bellman equation, incurs only a small computational overhead.